ES014A

PI: Shen
Title: 0106+013

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: A/dual pol
Obs time (UT): EU 15:00(19/02)-19:00(19/02)
Stations: EbHhJbMcNtOnShTrUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (GMaccaferri at Wed Feb 19 13:30:43 EST 1997) Success
All done.

HartRAO (Jonathan Quick at Wed Feb 19 14:09:01 EST 1997) Minor failures
Read heads failing - write heads still seem OK Only LCP available at HartRAO as yet

Effelsberg (dgraham at Thu Feb 20 02:54:51 EST 1997) Success
high wind but all observed

Westerbork (Tony Foley at Thu Feb 20 05:26:34 EST 1997) Success
No known problems

Onsala (A. Polatidis at Sun Feb 23 08:33:20 EST 1997) Success
No known problems. Note that Onsala records only LCP at 6 cm.

Jodrell (Alastair Gunn at Wed Mar 5 07:25:19 EST 1997) Failure
Did not observe due to severe winds all day.

Calibration (Fredrik T Rantakyro at Thu May 8 11:33:10 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name es14a.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Mon Mar 9 11:02:39 MET 1998) Success
Correlation completed. No problems found.